Determine the total​ first-year cost of car ownership for Milagros. She just purchased a vehicle valued for ​$21 comma 55021,550 with the following​ costs:

Auto Loan

​: Amount,

​$21,550​;

​Duration, 4​ years; APR,

8.25

percent

Property Taxes

​:

1.80

percent of vehicle​ value/year

Sales Taxes

​:

4.31

percent of the sales price

Title and Tags

​:

​$32.81​/year

Insurance

​:

​$2,247.58​/year

Maintenance and Usage

​:

​$1.52​/year
Answer #1

I have answered the question below

Please up vote for the same and thanks!!!

Do reach out in the comments for any queries

NOTE: As given in the question I have calculated the maintanence and usage as 1.52/yr. Please check the same and rectify the changes if needed as it is low

Answer:

APR = 8.25%

Vehicle purchase interest = 0.0825 * 21,550 = 1777.875

Property tax = 1.8% * 21,550 = 387.9

Sales Tax = 4.31% * 21,550 = 928.805

Other costs = 32.81 + 2247.58+1.52 = 2281.91

TCO = 2281.91 + 928.805 + 387.9+1777.875 = 5376.49